# A Unified Multi-Wavelength Data Analysis Workflow with gammapy

### v1

@@ Appendix D Systematic Uncertainties in Liverpool Telescope Data
When this component together with the source spectrum is fitted to the data, it causes a deformation of the assumed spectrum (e.g., a log-parabola), enabling the model to better align with the actual measurements (excess counts) by subtly adjusting the shape of the spectrum (see 13(b), light gray curve and black diamond markers). By removing this component, we can recover the underlying estimate of the true spectrum (dark gray) ,free from the influence of systematic errors. For clarity, in this figure we stacked together observations using the same filter, which in gammapy it is technically implemented in the stack_reduce method by co-adding the counts, background and exposures for datasets with compatible axes. The equivalent stacking is achieved for the native (classical) analysis by performing a weighted average of the flux point values, where the weights are set to the inverse of the square of the statistical errors of the points.
This correction effect is also illustrated in Figure 14(b), where arrows illustrate how the flux points would shift if the systematic components were excluded from the spectral model. The comparison between the two sets of spectra and flux points, one including the systematic component and one corrected for it, is shown in Figure 13(b).
Appendix E Second night analysis (MJD60384)
@@
The body of this manuscript primarily discusses the methodology used to create gammapy-compatible datasets, with data from the first night of NuSTAR observations (March 4th, 2024, MJD 60373) presented as an example of the reconstruction achievable using our analysis framework. In this appendix, we briefly report the results of applying the same analysis to the second night of NuSTAR observations (March 15th, 2024, MJD 60384) for which the temporal coverage of the instruments is shown in Figure 15.
Figure 15: Temporal coverage of the first observation night with the different participating instruments.
To keep this section short and focused on the final results, we focus on the two main figures from Section 4.4. The first figure is the best-fit models, shown in Figure 17, using the same sum of two power-laws with exponential cut-offs as in Figure 9. The corresponding corner plot, representing the error ellipses for each pair of free model parameters for the source emission component, is displayed in Figure 18. Compared to Figure 10, we observe a similar degree of degeneracy between the parameters for the low-energy (LE) component, and overall, there is a good agreement between the two fitting techniques for all parameters. However, in this case, the forward folding fit appears to better constrain the inverse of the cutoff energy $\lambda$ of the high-energy (HE) component, which is reflected in the figure as a more oblate ellipse. Regarding the parameters, the HE component seems to exhibit a slightly harder spectrum than on the first night.
We note however that there is some tension between Swift-UVOT and Liverpool IO:O, and between Swift-XRT and NuSTAR. We considered two possible causes of this. The first one is related to the observing times themselves in combination with intra-night source variability, as XRT observations are concentrated around the second half of NuSTAR only. This however is not supported by the relatively stable spectra in the Fermi-LAT, NuSTAR and Liverpool bands between the two nights. A more likely cause seem to be a degraded pointing of Swift satellite after one IRO (gyroscope) became noisy in Summer 2023 and then again in Spring 2024, leading to an increase chance of star tracker “loss of lock” after long slews. This is in fact hinted in our data from the Swift-XRT, whose integrated PC-mode image show a larger PSF and slightly larger field noise than that of the first night. And it is even more evident from the single UVOT exposure in filter UM2, where OP 313 image is far from round, and even not fully contained in the predefined source region size, as seen in Figure 16. While the effect should be less severe in XRT than in UVOT due to the larger intrinsic PSF, we believe this possible loss of lock in the star tracker during our observations may have led to a significant underestimation of the source flux in both instruments.
Figure 16: Swift-UVOT image with filter UM2 on the night of MJD60384. Blue circle indicate the standard ON integrated region, while the orange circles indicate the annulus for the background extraction.
@@
Figure 18: Corner plot showing the error ellipses calculated with a bootstrap method for the best-fit (free) parameters of the flux-point based fit and the full forward-folding joint fit. The model used for the LE is strongly degenerated because of the limited amount of available data to constrain that component.

### v2

@@ Appendix E Systematic Uncertainties in Liverpool Telescope Data
When this component together with the source spectrum is fitted to the data, it causes a deformation of the assumed spectrum (e.g., a log-parabola), enabling the model to better align with the actual measurements (excess counts) by subtly adjusting the shape of the spectrum (see Figure 13(b), light gray curve and black diamond markers). By removing this component, we can recover the underlying estimate of the true spectrum (dark gray), free from the influence of systematic errors. For clarity, in this figure we stacked together observations using the same filter, which in gammapy it is technically implemented in the stack_reduce method by co-adding the counts, background, and exposures for datasets with compatible axes. The equivalent stacking is achieved for the native (classical) analysis by performing a weighted average of the flux points, where the weights are set to the inverse of the square of the statistical uncertainties of the points.
Appendix F Second night analysis (MJD60384)
@@
The body of this manuscript primarily discusses the methodology used to create gammapy-compatible datasets, with data from the first night of NuSTAR observations (March 4th, 2024, MJD 60373) presented as an example of the reconstruction achievable using our analysis framework. In this appendix, we briefly report the results of applying the same analysis to the second night of NuSTAR observations (March 15th, 2024, MJD 60384). The corresponding temporal coverage is still consistent across all instruments, as seen in Figure 15.
To keep this section short and focused on the final results, we focus on the two main figures from Section 4.4. The first figure is the best-fit models, shown in Figure 16, using the same sum of two power-laws with exponential cut-offs as in Figure 9. We note that Swift-UVOT has a single observation with filter UM2, but more importantly, as discussed in section 3.4, Swift was affected by a noticeable degradation in the pointing accuracy caused by noise in one gyroscope (Cenko 2024a, b). This is in fact hinted in our Swift-XRT data, whose integrated PC-mode image shows a larger PSF than usual, and it is readily visible in UVOT, where OP 313 images are elongated and distorted, requiring an abnormally large source integration region of $25"$ radius to fully contain the signal. The standard $5"$ radius yields an underestimation of the source flux by a factor $\sim 2$.
The corresponding corner plot, representing the error ellipses for each pair of free model parameters for the source emission component, is displayed in Figure 17. Compared to Figure 10, we observe a similar degree of degeneracy between the parameters for the low-energy (LE) component. Moreover, despite our efforts to enlarge the source extraction region in both Swift-XRT and Swift-UVOT to accommodate their poorer PSF, a notable discrepancy persists between Swift-XRT and the spectral reconstruction from NuSTAR and Fermi-LAT. This inconsistency leads to less stable fitting process and different best-fit solutions for the low-energy component between the flux point fitting analysis and the forward folding technique.
It is worth noting that both the significant model degeneracies and fit convergence issues are common challenges faced when working with large and diverse datasets. However, these issues could potentially be mitigated by adopting a modern, fully Bayesian nested sampling inference engine such as UltraNest (Buchner 2021), which efficiently explores the entire prior parameter space to build robust posterior probability distributions for model parameters, even if the distribution is multi-modal. With the full dataset exported into a standard framework and the construction of a full forward folding likelihood, we are already halfway to exploring this approach.
Figure 15: Temporal coverage of the second observation night.
@@
Figure 17: Corner plot showing the error ellipses calculated with a bootstrap method for the best-fit (free) parameters of the flux-point based fit and the full forward-folding joint fit. The model used for the LE is highly degenerate because of the limited amount of available data to constrain that component.